OpenStudy (anonymous):

FAN AND MEDAL! What is the slope of the line define by the equations 8x+2y=5

OpenStudy (anonymous):

The answer is -4 All you do is solve for y so set up 8x+2y=5 subtract 8x from both sides 2y=-8x+5 no divide both sides by 2 y=-4x+5/2 m=-4 there you go And to remember it just write y=mx+b over and over again (m is the slope and b is the y intercept(y intercept is where the line crosses the y axis))

OpenStudy (anonymous):

thanks @Andyed52

OpenStudy (anonymous):

This belongs in the math subject not the english subject. To find the math subject, click Find More Subjects, and click the math subject. Also, close this since it's in the wrong spot. Thanks!

OpenStudy (anonymous):

oops sorry I thought I was in the math

OpenStudy (anonymous):

It's ok. Just close it.